The mean result from some experiment with 50 observations was 1.5 and the standard deviation was 1. Let LaTeX: \muμ denote the true mean result. We wish to to test LaTeX: H_0\colon \mu=1 H 0 : μ = 1 vs LaTeX: H_a \colon \mu \neq 1 H a : μ ≠ 1 . The p-value is

Group of answer choices

LaTeX: 2 \times P(t_{49} \leq 3.5)2×P(t49≤3.5)

LaTeX: 2 \times P(t_{49} \geq \left|3.5\right|)2×P(t49≥|3.5|)

LaTeX: 2 \times P(t_{49} \leq -3.5)2×P(t49≤−3.5)

LaTeX: 2 \times P(t_{49} \leq \left|3.5\right|)
Answer #1

Solution :

This is the two tailed test .

Test statistic = t

= (\bar x - \mu ) / s / \sqrt n

= (1.5 - 1) / 1 / \sqrt 50

Test statistic = 3.5

df = 50 - 1 = 49

The p-value is :

2 x P(t49 > |3.5|)